OpenStudy (zolock):

What are the factors of 1008?

OpenStudy (polaris_s0i):

easiest way to figure this out is to prime factorize: 1008 / 2 = 504 504 / 2 = 252 252 / 2 = 126 126 / 2 = 63 63 / 7 = 9 9 / 3 = 3 so the prime factorization becomes: \[2^4*3^2*7\] then you need to use combinations of those numbers to find the factors: \[ 2 * 504 = 2 * (2^3*3^2*7)\] \[3 * 336 = 3 * (2^4*3*7)\] \[4 * 252 = 2^2 * (2^2*3^2*7)\] \[6 * 168 = (2*3) * (2^3*3*7)\] \[7 * 144 = 7 * (2^4*3^2)\] \[8 * 126 = 2^3 * (2*3^2*7)\] \[9 * 112 = 3^2 * (2^4*7)\] \[12 * 84 = (2^2*3) * (2^2*3*7)\] \[14 * 72 = (2*7) * (2^3*3^2)\] \[16 * 63 = 2^4 * (3^2*7)\] \[18 * 56 = (2*3^2) * (2^3*7)\] \[21 * 48 = (3*7) * (2^4*3)\] \[24 * 42 = (2^3*3) * (2*3*7)\] \[28 * 36 = (2^2*7) * (2^2*3^2)\] you know you are done when you can't find any more combinations below the squareroot of 1008, which is about 31.74.